OpenStudy (anonymous):

Solve the equation for x if ax + bx = c x = c/(a + b) x = (a + b)/c x = c - a - b x = c/(ab)

OpenStudy (anonymous):

i need a awnser

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Oh, okay.

OpenStudy (igreen):

\(ax + bx = c\) \(a + b = c\) \(x = \dfrac{c}{a+b}\) So your answer is A.

OpenStudy (anonymous):

thank you if i need you for more will you be willing to help.

OpenStudy (igreen):

Of course. Just tag me.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The sum of Fred's age and ten is the same as twice his age minus 4. How old is Fred? 7 years old 6 years old 14 years old 18 years old @iGreen

OpenStudy (igreen):

The answer is C. Twice his age minus 4 = His age + 10 Just plug in 14 for his age.. Twice 14 minus 4 = 14 + 10 Twice 14 = 28 28 minus 4 = 14 + 10 24 = 14 + 10 24 = 24 So the answer is C!

OpenStudy (anonymous):

Solve the equation for r if |r| - 12 = 23 r = 11 r = -11, 35 r = -35, 35 r = 11, -11

OpenStudy (anonymous):

is it b are c @iGreen

OpenStudy (anonymous):

huh

OpenStudy (igreen):

C

OpenStudy (igreen):

-35 and 35 are both solutions..

OpenStudy (anonymous):

Olga is trying to figure out the length of a pool at a hotel that is to be built. A one inch door in the model will become a 3 foot door in reality. If the length of the pool is 7 inches on the model, how long will it be when it is built? 21 feet 10 feet 23 feet 13 feet

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

@iGreen